OpenStudy (anonymous):

A car dealer offers you two deals on a car that costs $16,000. Please calculate the monthly payment, given these two payment options the car dealer is offering. Payment Option 1: You can finance the car for 60 months with no interest if you make a $3,000 down payment. Payment Option 2: You can finance the car for 72 months (6 years) with 1% simple annual interest and no down payment. (Hint: To calculate simple annual interest, use the formula Interest = Principal * Rate * Time (in years). Add the amount of interest to the price of the car.) Which monthly payment amount is lower? Please exp

OpenStudy (anonymous):

Payment option 1 : 16,000 - 3000(down payment) = 13,000 13,000 for 60 months = 13,000/60 = 216.67 per month Payment option 2 : I = P * R * T I = 16,000 * .01 * 6 I = 960 16,000 + 960 = 16,960 16,960/72 = 235.56 per month So we have : payment option 1 : 216.67 per month for 60 months payment option 2 : 235.56 per month for 72 months what do you think the answer is ?

OpenStudy (anonymous):

1